OpenStudy (anonymous):

Factorization help please! original equation:xsquared-ysquared + 2zx +2yz+27 -2z -1 The furthest step i got to: xsquared + (2z)x -[(y-1){y-(2z + 1)] What do you do now?

OpenStudy (mr.math):

There are more than one factorization for such an expression.

OpenStudy (mr.math):

\[x^2-y^2+2zx +2yz+27 -2z -1=(x+y)(x-y)+2z(x+y)+26-2z\] \[=(x-y)(x+y+2z)+2(12-z)\]

OpenStudy (anonymous):

(x+y)(x+2z-y)-2z+26

OpenStudy (mr.math):

Oh, I've got a better factorization.

OpenStudy (anonymous):

Mr.Math its 13-z at last i think

OpenStudy (anonymous):

I messed up the furthest step I got to I got : x squared+ (2z)x- [y squared - (2z+2)y +(2z+1) ] They just want factorization, no 'answers'

OpenStudy (mr.math):

Yeah, 13-z.

OpenStudy (anonymous):

\[(x+y)(x+2z-y)-2(z-13)\]

OpenStudy (anonymous):

And it is x+y which has to be taken common over there See the above post I gave the answer to your thought

OpenStudy (anonymous):

oh I got something completely different (x+y-1) (x-y+2z+1)

OpenStudy (anonymous):

Hopex its perfect dude U can go for it

OpenStudy (mr.math):

I don't think that's right. @hopex

OpenStudy (anonymous):

ohhh sry sry wait wait u will get something to be added at last some constant

OpenStudy (anonymous):

oh wait the original equation is actually : xsquared - y squared + 2zx+2yz+2y-2z-1 Messed that up sorry!

OpenStudy (anonymous):

ohhh dude see that first buddy i thought its z at first but I thought it is correct as given by u

OpenStudy (anonymous):

So I got to the step xsquared +(2z)x -[ysquared - (2z +2)y + (2z+1)] but how does that factor into xsquared +(2z)x - (y -1) [y-(2z +1)] I don't get that step^